# Bushfires and Balance: Proactive versus Reactive Policies in Prescribed Burning

### v1

@@ Abstract.
We introduce a model to describe the interplay between prescribed burning and bushfires based on a system of ordinary differential equations.
We show that the system possesses a unique steady-state, whose stability depends on the policy governing prescribed burning. Specifically, a Reactive Policy in which prescribed burning activities are increased in response to bushfire events risks to produce instability and requires an additional feedback control, while a Proactive Policy in which moderate prescribed burning replaces natural bushfires may be more effective to maintain a balanced state. A high regeneration factor is also helpful to ensuring the stability of the equilibrium state.
The results presented here should not be regarded as a definitive solution to the issue of prescribed burning, as numerous specific factors must be considered in any comprehensive policymaking approach. Nonetheless, the stability analysis developed in this paper suggests that policies driven primarily by reactive responses to bushfire events may be emotionally charged and prone to instability. These reactive policies would benefit from incorporating feedback control mechanisms. In contrast, strategies based on proactive prescribed burning in moderate quantities, focused on prevention rather than reaction, are likely to be more effective in maintaining long-term ecological stability.
@@ 1. Introduction
We will show that the model possesses a unique fixed point. The stability of this equilibrium depends on the policy adopted for the prescribed burning. Specifically, we will show that when prescribed burning activities are incremented as a result of bushfire events, the equilibrium is unstable. Conversely, when the (moderate) prescribed burning replaces natural bushfires, the equilibrium is stable.
@@ 2.1. Modeling bushfires and prescribed burning
We suppose that the dynamics of this system is regulated by the following system of three ordinary differential equations. The first equation encodes the change in available land, which decreases due to bushfires and controlled burns and increases thanks to a regeneration factor. Specifically, we suppose that the available land decreases proportionally with respect to controlled burns, while the bushfires effect is linear with respect to both the available land and the bushfire intensity. As for the regeneration effect, we assume that it is proportional to the amount of burnt land (i.e., normalising to $1$ the total amount of land, to $1-a$). These considerations lead to the equation
$\dot{a}=-\alpha fa-\beta b+\gamma(1-a),$
@@
This equilibrium is already of practical interest. The relation between $a_{\star}$ and $f_{\star}$, that is,  $f_{\star}=\frac{\zeta a_{\star}}{\eta}$, quantifies the optimal bushfire intensity threshold in terms of the environmental land. The third relation in (2.4), being equivalent to $b_{\star}=\gamma-\frac{a_{\star}(\gamma\eta+\alpha\zeta a_{\star})}{\beta\eta}$, determines the optimal preventive burning given the environmental land.
For consistency, one expects the available area at equilibrium $a_{\star}$ to be bounded from above by the area of the full region, which has been normalised to $1$. This implies
@@
Regarding the units of measurement in our model, in addition to traditional land area metrics (e.g., acres), we mention that wildfire frequency can be expressed in acres per year (total area affected over time), years (average fire return interval at a location), or fires per unit area per year (a proxy for regional fire frequency). Similarly, wildfire intensity can be measured in megawatts, representing the total power emitted as thermal radiation, often assessed via satellite data. However, we emphasise that the key advantage of our model is its applicability as a conceptual framework, based on the overall structure of the problem more than on specific details, providing policymakers with early warnings about potential structural instabilities in policies set in place without a long-term plan.
@@ 2.2. Reactive Policy: Increasing prescribed burning after bushfires
Figure 1. Number of prescribed burns and wildfires in different areas of Western Australia, 2019–2022. Source: Brodie James Loller, 2022, %****␣unstable250221.tex␣Line␣325␣****https://storymaps.arcgis.com/stories/6428d63b54e4406eadb62e633f5f41fb.
@@
The equilibrium in (2.4) is linearly unstable when $\theta\in(0,+\infty)$, since the Jacobian matrix associated with the system in (2.3) and evaluated at (2.4) always possesses at least one real, positive eigenvalue.
More precisely, depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, this matrix can present:
  • three real eigenvalues, of which one is positive and two are negative; or
@@
There exist parameters that realise each of the cases above.
@@
For practical purposes, the statement of Theorem 2.1, highlighting the instability of the only physical equilibrium of the system, could be, in principle, interpreted as a negative result, hinting at significant concrete difficulties in managing bushfires merely on the basis of preventive burns. However, the quantitative analysis and the eigenvalue characterisation in Theorem 2.1 allows one to efficiently utilise an “augmentation method” or “dynamic extension”: more specifically, one can introduce an additional ordinary differential equation to the system to stabilise the equilibrium. This additional feedback mechanism influences the dynamics and provides a stabilising damping effect:
@@
While the statement proposed in Theorem 2.2 may sound vague and non-specific, the structure provided in Theorem 2.1 actually permits an explicit construction of this feedback mechanism, based on a partial diagonalisation of the Jacobian matrix and the coupling with a suitable, linear reaction term: in this sense, the proof of Theorem 2.2 is more interesting than its statement, since it exhibits a “constructive” procedure to stabilise the fixed point. Indeed, from a mathematical perspective, this feedback mechanism penalises the positive eigenvalue direction, making its identification useful for wildfire management. Furthermore, recognising the potential risk associated with instability can help justify additional preventive measures to regulate fuel availability.
@@ 2.3. Proactive Policy: Increasing (with moderation) prescribed burning in the absence of bushfires
Interestingly, another method to stabilise the equilibrium in (2.4) consists in modifying the strategy of the preventive burning. For example, one could perform less controlled burning in the presence of a highly intense natural fire activity. The ansatz in this case is that the severe bushfire events have already removed the available fuel from the soil, hence less preventive activity is necessary. Low bushfire activity leaves excessive fuel in the land which needs to be preventively removed via controlled burning.
@@
since the eigenvalues associated with the Jacobian matrix each have negative real parts.
More precisely, under assumption (2.7), depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, this matrix can present:
  • three real, negative eigenvalues,
  • one real, negative eigenvalue and two complex conjugate eigenvalues with negative real part.
There exist parameters that realise each of the cases above.
@@
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)<\beta\eta\vartheta f_{0},$
@@
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)=\beta\eta\vartheta f_{0},$
@@
Condition (2.7) is interesting since, being satisfied for small values of $\vartheta$, it suggests that moderate (rather than excessive) controlled burning could provide more stable outcomes. We can further determine the relevant factors in making decisions on controlled burning. In environments with low levels of natural regeneration, less prescribed burning is desirable, as the factor on the left-hand side of (2.7) will be lower. If the regenerative factor $\gamma$ is high, however, or the optimal bushfire intensity $f_{0}$ is high, then condition (2.7) is satisfied and more controlled burning remains permissible without sacrificing stability.
@@ 2.6. Organisation of the paper
The rest of this paper is organised as follows. The proofs of the above results are presented in Section 3, with the demonstration of some auxiliary algebraic results deferred to the appendix. Some conclusions will be summarised at the end of the paper.
@@ 3.1. Some preliminary algebraic observations
In the proof of Theorems 2.1 and 2.3, when establishing that the system may possess, depending on the paramenters, real eigenvalues as well as complex ones, the following observations will come in handy:
@@
$\lambda^{3}+B\lambda^{2}+C\lambda+E=0,$
@@
Since $P(0)=-D<0$ and $P(+\infty)=+\infty$, there exists at least one root $\lambda_{1}$ of $P$ lying in the interval $(0,+\infty)$, which proves that the equilibrium in (2.4) is linearly unstable.
@@
$\lambda^{3}+B\lambda^{2}+C\lambda-D=P(\lambda)=(\lambda-\lambda_{1})(\lambda-\lambda_{2})(\lambda-\lambda_{3})$
$\qquad=\lambda^{3}-(\lambda_{1}+\lambda_{2}+\lambda_{3})\lambda^{2}+(\lambda_{1}\lambda_{2}+\lambda_{1}\lambda_{3}+\lambda_{2}\lambda_{3})\lambda-\lambda_{1}\lambda_{2}\lambda_{3},$
@@
$-B=\lambda_{1}+\lambda_{2}+\lambda_{3}\qquad{\mbox{and}}\qquad D=\lambda_{1}\lambda_{2}\lambda_{3}.$
@@
$U=\left(\begin{matrix}\lambda_{1}&*&*\\ 0&\lambda_{2}&*\\ 0&0&\lambda_{3}\end{matrix}\right).$
@@
$x_{\star}:=Qy_{\star}\qquad{\mbox{and}}\qquad X(x):=QY(Q^{-1}x).$
@@
To stabilise the equilibrium $x_{\star}$ (or equivalently, up to a linear mapping, to stabilise the original equilibrium $y_{\star}$), we propose the following system of ordinary differential equations which includes now a feedback function $\omega:\mathbb{R}\to\mathbb{R}$:
@@
$J_{\star\star}:=\left(\begin{matrix}\lambda_{1}&*&*&-\tau\lambda_{1}\\ 0&\lambda_{2}&*&0\\ 0&0&\lambda_{3}&0\\ \sigma&0&0&-\sigma\end{matrix}\right).$
@@
$\lambda_{2},\quad\lambda_{3},\quad\frac{\lambda_{1}-\sigma+\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2},\quad{\mbox{and}}\quad\frac{\lambda_{1}-\sigma-\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2}.$
@@
$P(\lambda):=\lambda^{3}+\left(\alpha f_{0}+\gamma+\eta\right)\lambda^{2}+\left(2\alpha f_{0}\eta+\gamma\eta\right)\lambda+\beta\eta\vartheta f_{0}.$
@@
$\begin{split}&P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+D,\\ &{\mbox{with\leavevmode\nobreak\ $B$, $C$, $D\in(0,+\infty)$.}}\end{split}$
@@
$P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+BC=(\lambda+B)(\lambda^{2}+C),$
@@
$\dot{b}(0)-\dot{a}(0)=\theta(f(0)-f_{0})a(0)+\alpha f(0)a(0)+\beta b(0)-\gamma(1-a(0))$
$\quad=\big{(}\alpha f(0)+\beta+\gamma\big{)}a(0)-\gamma>0.$
@@
Hence, if, for the sake of contradiction, the claim in (A.1) were false, then $P$ would have three real roots, say $-a$, $-b$ and $-c$, with $a$, $b$, $c\in(0,+\infty)$.
@@
$P(\lambda)=(\lambda+a)(\lambda+b)(\lambda+c)=\lambda^{3}+(a+b+c)\lambda^{2}+(ab+ac+bc)\lambda+abc$
@@
To this end, from the above analysis we know that the three roots of $P$ have the form $-a$, $\alpha+i\beta$ and $\alpha-i\beta$, with $a\in(0,+\infty)$, $\alpha\in\mathbb{R}$, and $\beta\in\mathbb{R}\setminus\{0\}$, and our goal is to show that
@@
$P(\lambda)$ $=$ $(\lambda+a)(\lambda-\alpha-i\beta)(\lambda-\alpha+i\beta)$
$=$ $\lambda^{3}+(a-2\alpha)\lambda^{2}+(\alpha^{2}-2a\alpha+\beta^{2})\lambda+a(\alpha^{2}+\beta^{2}).$
@@
$B=a-2\alpha,\quad C=\alpha^{2}-2a\alpha+\beta^{2}\quad{\mbox{and}}\quad D=a(\alpha^{2}+\beta^{2}).$
@@
00 0 $<$ $D-BC$
$=$ $a(\alpha^{2}+\beta^{2})-(a-2\alpha)(\alpha^{2}-2a\alpha+\beta^{2})$
$=$ $2\alpha(a^{2}-2a\alpha+\alpha^{2}+\beta^{2})$
$=$ $2\alpha\big{(}(a-\alpha)^{2}+\beta^{2}\big{)},$
@@
$B=(1+2c)\alpha,\qquad C=c(2+c)\alpha^{2}\qquad{\mbox{and}}\qquad E=-c\alpha\beta\theta,$

### v2

@@ Abstract.
We introduce a new mathematical model to explore the dynamic relationship between prescribed burning and bushfire occurrence, formulated as a system of ordinary differential equations. The model admits a unique steady-state, and its stability is shown to depend critically on the policy framework governing prescribed burning. In particular, reactive policies, where prescribed burning is increased in response to bushfire events, can lead to system instability unless augmented with appropriate feedback control. Conversely, proactive policies that implement moderate, consistent prescribed burning are more effective at stabilizing the system and mitigating the frequency of bushfires. Additionally, higher vegetation regeneration rates contribute positively to equilibrium stability. While this model does not capture the full complexity of fire management, it highlights the potential risks of emotionally driven, reactive responses and underscores the value of preventive, stability-focused strategies. To the best of our knowledge, this is the first mathematical analysis of preventive burning strategies and their long-term impact on bushfire mitigation.
@@ 1. Introduction
The scientific literature on wildfires is extensive, particularly with regard to mathematical models and software tools developed to simulate fire front propagation (see e.g. [BEER, MR4772545, SPARK] and references therein).
In parallel, a distinct body of research has examined the social dimensions of bushfire prediction and the role of bushfire simulators in policy-making and operational decision-making (see e.g. [NEA1qeA]).
To the best of our knowledge, however, this work is the first to present a mathematical analysis of preventive burning strategies and to investigate how different long-term management approaches and evolving trends influence the stable mitigation of bushfire events over extended time horizons.
@@
We will show that the model possesses a unique fixed point. The stability [^1] The stability of an equilibrium in a model of physical relevance is of paramount importance for practical applications: according to [MR2144536, page 175], “one cannot usually pinpoint positions exactly, but only approximately, so an equilibrium must be stable to be physically meaningful”. of this equilibrium depends on the policy adopted for the prescribed burning. Specifically, we will show that when prescribed burning activities are incremented as a result of bushfire events, the equilibrium is unstable. Conversely, when the (moderate) prescribed burning replaces natural bushfires, the equilibrium is stable.
@@ 2.1. Modeling bushfires and prescribed burning
We suppose that the dynamics of this system is regulated by the following system of three ordinary differential equations. The first equation encodes the change in available land, which decreases due to bushfires and controlled burns and increases thanks to a regeneration factor. Specifically, we suppose that the available land decreases proportionally with respect to controlled burns, while the bushfires effect is linear with respect to both the available land and the bushfire intensity. As for the regeneration effect, we assume that it is proportional to the amount of burnt land (i.e., normalising to $1$ the total amount of land, to 1a1-a). These considerations lead to the equation
(2.1) $\dot{a}=-\alpha fa-\beta b+\gamma(1-a),$
@@
The term $\eta f$ represents the decay of bushfire activity in the absence of available land $a$, which may result from factors such as fuel depletion or suppression efforts by fire brigades.
@@
As is customary in the analysis of differential equations (see e.g. [MR1908418, page 3]), our mathematical framework allows the involved variables to take negative values, in line with the analytic formulation of the system. However, in the context of our application, the quantities of interest, namely bushfire frequency, available land, and controlled burns, are assumed to be nonnegative, reflecting their real-world interpretation.
@@
This equilibrium is already of practical interest. The relation between $a_{\star}$ and $f_{\star}$, that is,  $f_{\star}=\frac{\zeta a_{\star}}{\eta}$, quantifies the optimal bushfire intensity threshold in terms of the environmental land. The third relation in (2.5), being equivalent to $b_{\star}=\frac{\gamma}{\beta}-\frac{a_{\star}(\gamma\eta+\alpha\zeta a_{\star})}{\beta\eta}$, determines the optimal preventive burning given the environmental land.
For consistency, one expects the available area at equilibrium $a_{\star}$ to be bounded from above by the area of the full region, which has been normalized to $1$. This implies
@@
When this condition is violated, the prescribed burning strategy becomes physically unfeasible, resulting either in the complete consumption of the available area through burning or in the absence of any preventive burning activity. In this context, condition (2.7) is particularly informative, as it establishes quantitative thresholds that determine the viability of prescribed burning strategies. Specifically, the parameter $\zeta$, which reflects the flammability of the available fuel, must lie within a precise range: namely, it must be sufficiently large to enable effective fuel reduction, yet not so large as to induce uncontrollable burning. These thresholds are explicitly captured by the double inequality in (2.7), offering practical insight into the constraints under which prescribed burning can be successfully implemented.
@@
Regarding the units of measurement in our model, in addition to traditional land area metrics (e.g., acres), we mention that wildfire frequency can be expressed in acres per year (total area affected over time), years (average fire return interval at a location), or fires per unit area per year (a proxy for regional fire frequency). Similarly, wildfire intensity can be measured in megawatts, representing the total power emitted as thermal radiation, often assessed via satellite data (see Table 1). However, we emphasize that the key advantage of our model is its applicability as a conceptual framework, based on the overall structure of the problem more than on specific details, providing policymakers with early warnings about potential structural instabilities in policies set in place without a long-term plan.
State variables and parameters Physical dimensions Methods of measurement
Land area ($a$, $b$) Metric units (e.g., acres) Manual surveys, GPS devices
Wildfire frequency ($f$, $f_{0}$)
Acres per year (total area affected over time),
years (mean fire return interval), or
fires per unit area per year
Historical records, direct observation
Wildfire intensity (alternative $f$, $f_{0}$) Megawatts (thermal radiation power) Satellite remote sensing
Proportionality parameters Dimensionally consistent (unit-dependent) Estimated from empirical data
Table 1. State variables and parameters: physical dimensions and methods of measurement
@@ 2.2. Reactive Policy: Increasing prescribed burning after bushfires
Figure 1. Number of prescribed burns and wildfires in different areas of Western Australia, 2019–2022. Source: Brodie James Loller, 2022, https://storymaps.arcgis.com/stories/6428d63b54e4406eadb62e633f5f41fb.
@@
The equilibrium in (2.5) is linearly unstable when $\theta\in(0,+\infty)$.
More precisely, depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, the Jacobian matrix associated with the system in (2.4) and evaluated at (2.5) presents:
  • either three real eigenvalues, of which one is positive and two are negative; or
@@
There exist parameters that realize each of the cases above.
@@
For practical purposes, the statement of Theorem 2.1, highlighting the instability of the only physical equilibrium of the system, could be, in principle, interpreted as a negative result, hinting at significant concrete difficulties in managing bushfires merely on the basis of preventive burns. However, the quantitative analysis and the eigenvalue characterisation in Theorem 2.1 allows one to efficiently utilize an “augmentation method” or “dynamic extension”: more specifically, one can introduce an additional ordinary differential equation to the system to stabilize the equilibrium. This additional feedback mechanism influences the dynamics and provides a stabilising damping effect:
@@
While the statement proposed in Theorem 2.2 may sound vague and non-specific, the structure provided in Theorem 2.1 actually permits an explicit construction of this feedback mechanism, based on a partial diagonalisation of the Jacobian matrix and the coupling with a suitable, linear reaction term: in this sense, the proof of Theorem 2.2 is more interesting than its statement, since it exhibits a “constructive” procedure to stabilize the fixed point. Indeed, from a mathematical perspective, this feedback mechanism penalizes the positive eigenvalue direction, making its identification useful for wildfire management. Furthermore, recognising the potential risk associated with instability can help justify additional preventive measures to regulate fuel availability.
@@ 2.3. Proactive Policy: Increasing (with moderation) prescribed burning in the absence of bushfires
Interestingly, another method to stabilize the equilibrium in (2.5) consists in modifying the strategy of the preventive burning. For example, one could perform less controlled burning in the presence of a highly intense natural fire activity. The ansatz in this case is that the severe bushfire events have already removed the available fuel from the soil, hence less preventive activity is necessary. Low bushfire activity leaves excessive fuel in the land which needs to be preventively removed via controlled burning.
@@
(2.9) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)>\beta\eta\vartheta f_{0}.$
More precisely, under assumption (2.9), depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, the Jacobian matrix associated with the system in (2.4) and evaluated at (2.5) presents:
  • either three real, negative eigenvalues,
  • red one real, negative eigenvalue and two complex conjugate eigenvalues with negative real part.
There exist parameters that realize each of the cases above.
@@
(2.11) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)=\beta\eta\vartheta f_{0},$
@@
Condition (2.9) is interesting since, being satisfied for small values of $\vartheta$, it suggests that moderate (rather than excessive) controlled burning could provide more stable outcomes. We can further determine the relevant factors in making decisions on controlled burning. In environments with low levels of natural regeneration, less oscillations in the prescribed burning program (obtained by reducing $\vartheta$) are desirable, as the factor on the left-hand side of (2.9) will be lower. If the regenerative factor $\gamma$ is high, however, or the optimal bushfire intensity $f_{0}$ is high, then condition (2.9) is satisfied and higher oscillations in the controlled burning program remain permissible without sacrificing stability.
The transition to instability, also passing through the critical case in (2.11), is visualized in Figures 4 and 5.
Refer to captionRefer to caption
Figure 4. Sketch of the streamlines of (2.4). We adopt here the choice of parameters $\alpha:=\beta:=\gamma:=\eta:=f_{0}:=1$ and $\zeta:=2.5$, which satisfies (2.6) and (2.7), and $\vartheta\in\{0.1,\,3,\,20\}$. These values of $\vartheta$ exemplify, respectively, (2.9), (2.11), and (2.10). The images confirm the transition towards instability.
@@ 2.4. Further observations
We also emphasize that the results in Theorems 2.4 and 2.5 do not imply that the model presented here is unrealistic. On the contrary, Theorem 2.4 should be interpreted as a caution against burning too much land (potentially all the available area) through control burnings. Meanwhile, Theorem 2.5 serves as a consistency result, provided that the available region is not exhausted.
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 5. Trajectories of the streamlines of (2.4) with the choice of parameters $\alpha:=\beta:=\gamma:=\eta:=f_{0}:=1$, $\zeta:=2.5$, and $\vartheta\in\{0.1,\,3,\,20\}$. These values of $\vartheta$ exemplify, respectively, (2.9), (2.11), and (2.10). The initial points are 100 random perturbations of the equilibrium in (2.5). The maximal size of the perturbation is $\pm 0.0001$. The images confirm the transition towards instability.
@@ 2.6. Organisation of the paper
The rest of this paper is organized as follows. We give a sketch of the proofs of the main theorems in Section 3. The detailed proofs of these results are presented in Section 4, with the demonstration of some auxiliary algebraic results deferred to the appendix. Some conclusions will be summarized at the end of the paper.
3. Insights into the main proofs
One of the main focuses of this paper is the linear stability and instability analysis of the new model for prescribed burning that we have introduced. The methodology employed is classical, relying on the identification of equilibria and the spectral analysis of the linearized system. In particular, complex eigenvalues of the Jacobian matrix with positive real parts correspond to unstable modes: oscillatory patterns arise when the imaginary part is nonzero, while real, positive eigenvalues indicate instabilities along escaping directions. Conversely, eigenvalues with negative real parts correspond to stable configurations.
The spectral analysis in our case is nontrivial, as the characteristic polynomial of the Jacobian matrix is of third degree. While, in principle, such a polynomial can be explicitly factorized, the resulting expression would be cumbersome and practically intractable, especially due to the presence of several structural parameters. Therefore, to determine the stability or instability of the system, we develop and employ a number of bespoke algebraic arguments.
Specific arguments from the theory of ordinary differential equations will also be used to analyze the trajectories in particular cases of interest.
The mathematical methods employed are elementary in nature, requiring no advanced background. However, they are carefully tailored to suit the model under consideration and to highlight the connection between the theoretical findings and their practical implications for controlled burning planning.
@@ 4.1. Some preliminary algebraic observations
The spectral analysis of the linearized system is not standard, since explicit solutions, though available, would be practically difficult or impossible to manage. To circumvent this difficulty, we introduce here some auxiliary algebraic results. Indeed, in the proof of Theorems 2.1 and 2.3, when establishing that the system may possess, depending on the paramenters, real eigenvalues as well as complex ones, the following observations will come in handy:
@@
(4.4) $\lambda^{3}+B\lambda^{2}+C\lambda+E=0,$
@@
Since P(0)=D<0P(0)=-D<0 and $P(+\infty)=+\infty$, there exists at least one root $\lambda_{1}$ of $P$ lying in the interval $(0,+\infty)$, which proves that the equilibrium in (2.5) is linearly unstable.
@@
(4.9) $\begin{split}&\lambda^{3}+B\lambda^{2}+C\lambda-D=P(\lambda)=(\lambda-\lambda_{1})(\lambda-\lambda_{2})(\lambda-\lambda_{3})\\ &\qquad=\lambda^{3}-(\lambda_{1}+\lambda_{2}+\lambda_{3})\lambda^{2}+(\lambda_{1}\lambda_{2}+\lambda_{1}\lambda_{3}+\lambda_{2}\lambda_{3})\lambda-\lambda_{1}\lambda_{2}\lambda_{3},\end{split}$
@@
(4.10) $B=\lambda_{1}+\lambda_{2}+\lambda_{3}\qquad{\mbox{and}}\qquad D=\lambda_{1}\lambda_{2}\lambda_{3}.$
@@
(4.15) $U=\left(\begin{matrix}\lambda_{1}&*&*\\ 0&\lambda_{2}&*\\ 0&0&\lambda_{3}\end{matrix}\right).$
@@
(4.16) $x_{\star}:=Qy_{\star}\qquad{\mbox{and}}\qquad X(x):=QY(Q^{-1}x).$
@@
To stabilize the equilibrium $x_{\star}$ (or equivalently, up to a linear mapping, to stabilize the original equilibrium $y_{\star}$), we propose the following system of ordinary differential equations which includes now a feedback function $\omega:\mathbb{R}\to\mathbb{R}$:
@@
The practical implementation of feedback mechanisms in controlled burns is inherently complex, and necessarily so, given the structural instability of the underlying system. At the core of this approach lies a quantitative analysis of a derived variable, denoted $x_{1}$, which captures the aggregate effects of the original variables $(a,f,b)$, representing available fuel, bushfire intensity, and the extent of controlled burns, respectively. We remark that this quantity $x_{1}$ is not an abstract construct: in fact, once the original variables are measured, $x_{1}$ is obtained via a simple linear combination. The coefficients involved in this combination are explicitly determined through the above Schur triangulation process, making the formulation of the feedback mechanism fully transparent.
The primary objective of the feedback is to dampen the destabilizing influence of $x_{1}$, which is tracked through the evolution of the variable $\omega$ which quantifies the required adjustment to $x_{1}$ in order to restore stability to the system. Translating this adjustment back into the original variables provides an explicit prescription for modifying $(a,f,b)$ to stabilize bushfire dynamics.
However, this process also highlights a fundamental modeling challenge: in real-world bushfire management, policy interventions can directly influence controlled burns $b$ and, to some extent, the amount of available fuel (thus affecting the values of $a$) through land management practices. In contrast, bushfire intensity $f$ is largely governed by unpredictable natural factors and thus lies outside direct control. Despite this limitation, the feedback mechanism described offers a valuable starting point, namely an explicit tool for guiding interventions. While it may not fully account for all practical constraints, especially regarding the uncontrollable variable $f$, it serves as a first-order approximation that can be enhanced with heuristic and empirical strategies. Such an approach can be particularly effective when the system operates near an optimal fire activity level $f_{0}$, which possibly does not requires major adjustments in the variable $f$. Additionally, the explicit expression of the feedback control in certain parameter ranges may suggest indirect actions for policymakers to intervene in bushfire intensity, such as increasing wildfire surveillance, enlarging fire brigades, enhancing early warning systems, improving forest management practices, or allocating more resources to community preparedness and rapid response strategies.
@@
(4.20) $J_{\star\star}:=\left(\begin{matrix}\lambda_{1}&*&*&-\tau\lambda_{1}\\ 0&\lambda_{2}&*&0\\ 0&0&\lambda_{3}&0\\ \sigma&0&0&-\sigma\end{matrix}\right).$
@@
(4.21) $\lambda_{2},\quad\lambda_{3},\quad\frac{\lambda_{1}-\sigma+\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2},\quad{\mbox{and}}\quad\frac{\lambda_{1}-\sigma-\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2}.$
@@
(4.22) $P(\lambda):=\lambda^{3}+\left(\alpha f_{0}+\gamma+\eta\right)\lambda^{2}+\left(2\alpha f_{0}\eta+\gamma\eta\right)\lambda+\beta\eta\vartheta f_{0}.$
@@
(4.23) $\begin{split}&P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+D,\\ &{\mbox{with\leavevmode\nobreak\ $B$, $C$, $D\in(0,+\infty)$.}}\end{split}$
@@
(4.33) $P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+BC=(\lambda+B)(\lambda^{2}+C),$
@@
(4.34) $\begin{split}&\dot{b}(0)-\dot{a}(0)=\theta(f(0)-f_{0})a(0)+\alpha f(0)a(0)+\beta b(0)-\gamma(1-a(0))\\ &\quad=\big{(}\alpha f(0)+\beta+\gamma\big{)}a(0)-\gamma>0.\end{split}$
@@
Hence, if, for the sake of contradiction, the claim in (A.1) were false, then $P$ would have three real roots, say a-a, b-b and c-c, with $a$, $b$, $c\in(0,+\infty)$.
@@
(A.2) $P(\lambda)=(\lambda+a)(\lambda+b)(\lambda+c)=\lambda^{3}+(a+b+c)\lambda^{2}+(ab+ac+bc)\lambda+abc$
@@
To this end, from the above analysis we know that the three roots of $P$ have the form a-a, $\alpha+i\beta$ and $\alpha-i\beta$, with $a\in(0,+\infty)$, $\alpha\in\mathbb{R}$, and $\beta\in\mathbb{R}\setminus\{0\}$, and our goal is to show that
@@
(A.4) $\begin{split}P(\lambda)&=(\lambda+a)(\lambda-\alpha-i\beta)(\lambda-\alpha+i\beta)\\ &=\lambda^{3}+(a-2\alpha)\lambda^{2}+(\alpha^{2}-2a\alpha+\beta^{2})\lambda+a(\alpha^{2}+\beta^{2}).\end{split}$
@@
(A.5) $B=a-2\alpha,\quad C=\alpha^{2}-2a\alpha+\beta^{2}\quad{\mbox{and}}\quad D=a(\alpha^{2}+\beta^{2}).$
@@
(A.6) $\begin{split}0&<D-BC\\ &=a(\alpha^{2}+\beta^{2})-(a-2\alpha)(\alpha^{2}-2a\alpha+\beta^{2})\\ &=2\alpha(a^{2}-2a\alpha+\alpha^{2}+\beta^{2})\\ &=2\alpha\big{(}(a-\alpha)^{2}+\beta^{2}\big{)},\end{split}$
@@
Figure 6. Plot of $\triangle$ as a function of $\alpha$ when $f_{0}:=1$, $\gamma:=\eta:=c\alpha$, $\zeta:=(1+c)\alpha$, $\beta:=1$, and $c:=0.1$. Notice that $\triangle>0$ for large values of $\alpha$.
@@
(A.8) $B=(1+2c)\alpha,\qquad C=c(2+c)\alpha^{2}\qquad{\mbox{and}}\qquad E=-c\alpha\beta\theta,$